Can you please help me with this linear equation?
One number is 15 more than another. The sum of twice the larger and 3 times the smaller is 182. Find the numbers.
thanks

Answer by user_dude2008(1862)

"One number is 15 more than another." ----> y=x+15

"The sum of twice the larger and 3 times the smaller is 182" ----> 2y+3x=182


2y+3x=182


2(x+15)+3x=182


2x+30+3x=182


2x+3x=182-30

5x=152

x=30.4


y=x+15

y=30.4+15

y=45.4